NAOJ GW Elog Logbook 3.2

I installed Debian 11 on the Data Concentrator/Network Data Server/Frame Writer computer in TAMA. In 2951 we ran across a problem where we couldn't complete the software installation for the Debian 11 install, and for some reason couldn't install a boot loader properly. The problem was that the installation medium was running in BIOS mode as opposed to the more appropriate UEFI mode. Switching to UEFI installer upon startup made the installation go smoothly as for the frontend computer.
The following options were chosen during installation:
Language - English
Region - Japan
Locales - en_US.UTF-8
Keyboard - American English
hostname - DCCentral
domain name - mtk.nao.ac.jp (default)
Network interface - eno1 Intel Corporation I210 Gigabit Network Connection
root password - normal KAGRA controls password
ops user 1000 (same password as usual)
controls user 1001 (same password as usual)
package manager - deb.debian.org
HTTP proxy - blank
The following drive partitioning scheme was chosen:
Disk 1 - 1 TB Western Digital ATA - for operating system and user files
512 MB - EFI System Partition (ESP) - in the failing step last time, using the BIOS mode installer would not let us install this partition. Now it's fine.
128 GB swap space
Remaining (~870 GB) - ext4 / - perhaps there's no real need to separate into /home, /var, /tmp here, so the system can just figure that out by itself
Disk 2 - 24 TB Avago MegaRAID virtual drive (4x 7.6 TB disks with RAID5 redundancy) - for mass storage of channel data
24.0 TB - ext4 /data0 - used for frame writer. In KAGRA, there are two frame writer PCs k1fw0 /data0 and k1fw1 /data1 with about 25 TB storage each, for extra data redundancy. But we only have one server PC, so I just assume we are using one frame writer.
Afterwards, I added ops and controls to the sudoers file and installed cdssoft version 1.0.11 (was 1.0.9 when installing frontend) from the aligo caltech Debian Bullseye repository.
*I will comment later in a bit more detail about what some of these technical terms mean

Marc, Yuhang
We have to reload DDS2 config everytime to be able to lock.
Then we are looking at the GR_corr spectrum to investigate the driving matrix of the INPUT but we found out that the noise level is not stable at all (see figure 1).
Our reference level is in blue with injected line on INPUT_PIT_ex2 at 5 Hz and 5000 amplitude is blue curve of figure 1.
At other times, we can see noise increase up to 10 Hz or even at higher frequency.
This is true with or without line injected, with or without AA or pointing loop.
Same after restarting laser.
Not affected by closing AA or pointing loops.
After some time, the noise excess disappeared and we could start some measurements but it quickly reappeared again..
PIT_ex2 4500, GR_corr 7.52
4000, 6.70
3500, 6.47
3000, 5.97
2500, 5.38

Marc, Yuhang
We found that GR reflection was far from overlapping with injection.
This was due to huge offset on input and end because AA loop was still closed.
After opening the loop, we could see TEM00 flashes but we had to have about 100 offset in input and end pitch.
We could not lock the FC so we checked the GR power while changing the SHG temperature.
GR power after FI (mW): 41.7, 45, 35, 44, 45
SHG temperature: 3.085, 3.071, 3.06, 3.0665, 3.075
Finally, we tuned the SHG alignment and could recover more than 50 mW after the FI.
We could still not lock so we tried to tweak the servo parameters :
gr lock servo, attenuation = 0.2, gain = 8, both changed no diff
gr reflection power > 200 uW
Finally, after changing the lock threshold we locked on a sideband. Turning off the servo seems to change the sign of the error signal and we could lock the FC.
In the end, we changed sign on DDS2 by reloading the latest configuration.

I set IR beam power to 1.6mW, I checked vertical and horizontal AOI with the razor blades and got respectively -0.008 deg and -0.022 deg.
I tuned the QWP and HWP angle to minimize p polarization while injection s polarization.
I did the calibration and started measurement of #4 (was cleaned with first contact) with s polarization at the input.

Yuhang and Marc
We would like to have 'tilt driving' coupling to 'shift driving' as small as possible. To achieve this, we worked on finding optimal driving matrix for H1 and H3 for pitch today.
At the beginning, the driving matrix for pitch is: H1 0.794, H3 1. This was optimized in the past. However, we would like to optimize it again since we glued new magnets at the beginning of this year.
To check how much coupling we have, we sent 5Hz 5000 amplitude for INPUT_PIT_ex2. We checked coherence between GR_CORR and excitation, no coherence was found.
Then we checked when driving matrix is: H1 1, H3 1. We found no coherence even in this case. We checked the connection, we found later on that actually GR_CORR is not connected.
After putting back good connection of GR_Corr, we take data when H1 and H3 are both 1. We saved as reference as blue (REF0 and REF1).
H1 0.794, GR_corr 17.964 (REF2, REF3)
H1 1, GR_corr 30.475
H1 0.6, GR_corr 13.38
H1 0.5, GR_corr 15.82
H1 0.55, GR_corr 13.82
H1 0.65, GR_corr 13.09
H1 0.65, H2 0.02, GR_corr 12.63
H1 0.65, H2 0.05, GR_corr 13.61
H1 0.65, H2 0.01, GR_corr 12.30
H1 0.65, H2 -0.01, GR_corr 12.82
When we drive H1, we get GR_corr 87.08
When we drive H3, we get GR_corr 63.24
We found a ratio of 0.73. Then we did H1 0.73, we get GR_corr 10.71
Attached figure shows a comparison of GR_corr when H1 is 1 or 0.73.
We still have coupling as we see the number in GR_corr 10.71. We still would like to understand it better what is causing this number. One thing we will check is the coupling factor when different amplitude signal is sent to H1 and H3. We will also test this coupling number by sending excitation at a different frequency.

Aso, Marc, Matteo
We put marking on AZTEC #1 and #3 to indicate the orientation of the sapphire during shaping.
One large arrow on the top pointing towards the front surface during our measurement + 2 long lines on the side at the level of the holder.
We packed the samples inside their wooden crates with foam to prevent their movements during shipping.

Marc, Matteo
We rotated the sample by 90 degrees and fine tuned the rotation angle to minimize p polarization power at the readout.
Results are attached to this entry.
It seems that previous position was better so today we rotated the sample back to this position.
Actually, trying to minimize the p polarization after the sample, we rotated clockwise from injection part by an additional 2.54 deg.
From HWP, it seems that we are now at 0.6 deg from theta = 0 deg.

We noticed a mistake in the calibration step in the scatter measurement.
We assumed that the camera was receiving all of the light emitted from the integrating sphere, when in fact it was receiving only 0.43%.
This reduces the result for all scattered light to 0.43%.
In other words, the 7/22 result is 4~7%*4.3*10^(-3)= 1.7~3.0*10^(-2)% ,instead of 4~7%.
(The reason for the wide range of values is that the intensity of the scattered light has an angular dependence.)
This means that the scattered light is about 170~300ppm.

Following the solved issue of PSD saturation, I computed again calibration parameters (see figures 1 to 3).
Then, I computed polarization map with input polarization s (figure 4) and 15 deg (figure 5).
From these measurements, I could compute delta n and theta (figure 6) as well as polarization conversion losses (figure 7).
The 2 methods to estimate these losses agree well with each other and show losses at about 0.74 % !
We will fine tune the mirror rotation angle (seems we still have an offset about 2.37 deg or 4.65 arc length cord).

We calibrated IR camera sensitivity and analyze the scattering data.
According to our result, Aztec3 has 4~7% scattering for IR light.

Previous entries reported strange anti-correlation of s and p polarization power while not changing polarization.
Furthermore we had also s polarization power too constant while rotating the mirror.
Today I found out that the PSD were saturating despite neither the PSD displays nor the lockin-amplifier showing saturation.
This is because I forgot to reduce the laser power after checking the beam parameter with the razor blades.
The power is now decreased from 14 mW to 1.6 mW.
I redid the calibration measurement and started measurement with rotated sample and s polarization.

I found that green reflection for AA was misaligned a lot. I aligned steering mirrors for AA to make the green beam at the center of both QPDs. DC centering loop can be locked, but AA loop could not be locked. I checked AA setting and it was completely different from previous setting in elog2850 (Fig. 1). I brought it back to the previous setting and AA loop seems working now.
To check the IR alignment, I injected BAB to FC. I could not find BAB transmission, but I found BAB reflection from FC. I will maximize the BAB reflection and recover IR alignment soon.

With Homare Abe
Date: 2022/7/20
We performed a camera calibration measurement for the camera which we used to measure scattering of Aztec No3.
We used an integration sphere with a 1050nm LED (LED1050L2) input as a reference light, then we took pictures with conditions (distance, ISO, focal length etc) same as we took the Aztec sample.
We changed exposure time and took some data.
After the measurement, we converted the data to "count sum VS energy caught by the camera", which is attached here.
Now, we can convert picture data for Aztec No3 to energy or power.

Marc, Michael, Yuhang
We checked the p polarization power after sample at its center while injecting s polarization and rotating the sample.
Result is attached to this entry.
Note that here, the positive rotation is opposite of positive rotation of the HWP. It means that we expect to see the minimum of p polarization power around -41.5 deg.
We indeed find a minimum around this position and started new measurement with the mirror rotated by -46.5 degrees.
During this measurement we found out that we have 2 beams on each port of the PBS on the readout part (maybe due to birefringence induced beam splitting).

We measured AZTEC #1 polarization with 0 deg (or s polarization), 15 and 45 deg polarization angle at the injection (see figures 1 to 3).
From the measurements with 0 and 45 deg input polarization angle, I could compute delta n and theta (figure 4) and s to p polarization losses (figure 5).
From this last figure, we can see a good agreement of losses between the direct measurement with the computation from delta n and theta.
This was not the case for the measurement with 15 deg input polarization angle and investigation are on-going.
The bad news is that theta is not uniform inside the central area of the mirror. This means that there is no rotation of the sample that will decrease the s to p polarization losses inside this area.

We could see IR scattering light.(See fig.)
Scattered light was visible only in the center of substrate.
We measured again in another angle.
Scattered light was visible in same position.
So we believe that internal scattered light does not have a strong angular dependence.
We start calibration for IR measurement.
We will use 1050nm LED.
We will use an optical table outside clean room. Please don't touch these components.

To check the alignment of green from OPO at GRMC, I put a PD at GRMC reflection for green from OPO (Fig. 1). The cursor 2 at 11.2mV shows the offset. Fig. 1 would show the green alignment to OPO is fine.
Then I measured nonlinear gain with BAB maximum. After some alignment of green to OPO, nonlinear gain with 25 mW green reached 5. This is consistent with theoretical nonlinear gain of 5.1 with 80mW OPO threshold, but lower than 6.8 which is the value one year ago in elog2577. The current setting is as follows.
green power (mW) | 0 | 25 |
OPO temperature (kOhm) | 7.13 | 7.13 |
p pol PLL frequency (MHz) | 230 | 160 |
BAB maximum (mV) | 286 | 1420 |
Nonlinear gain | 1 | 5 |

I scanned BS alignment and green beam can be found in FC transmission. I aligned input mirror with picomotor so that the green reflection overlaps with the injection. After I aligned end mirror, green flash can be found and FC locked stably!

Marc, Matteo
We removed the AZTEC #5 from the translation stage and installed 2 razor blades.
They are at 120 mm (for the vertical cut) and 122 mm (for the horizontal cut) of the last steering mirror on the injection path.
We realigned the probe beam and reached 0.003 deg and 0.005 deg angle of incidence for vertical and horizontal respectively.
We realigned the PBS and our photodetectors, tuned the QWP and HWP.
S polarization is best for 3.1 deg of the HWP. The best p polarization is reached by moving the HWP angle by -45 deg.
We did the calibration as for #3.
A bit strange result was that s and p polarization power fluctuations were anti-correlated suggesting that the input polarization was fluctuating.
Touching a bit the laser fiber I could generate such anti-correlation. One possible explanation for this could be that the fiber is now touching/resting on the scattering black box foil which might be moving due to the air flow (?).
In any case, we could get some reasonable calibration factors but we will further investigate this issue.
Finally, we started polarization measurement with s polarization at the input.

Report about AZTEC #5 absorption.
Absorption is about 140 ppm/cm
Interestingly, it is the only sample with so visible crystalline structure defects (see yz map).
It also had largest scattering area (visible with our probe laser) compared to the other samples.